16. Finding the WACC. Hankins Corporation has 5.4 million shares of common stock outstanding, 290,000 shares of 5.6 percent pPlease show work not done on excel. Thanks
Answer #1

a

MV of equity=Price of equity*number of shares outstanding
MV of equity=64*5400000
=345600000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*125000*1.09
=136250000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=103*290000
=29870000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=345600000+136250000+29870000
=511720000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 345600000/511720000
W(E)=0.6754
Weight of debt = MV of Bond/MV of firm
Weight of debt = 136250000/511720000
W(D)=0.2663
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 29870000/511720000
W(PE)=0.0584

b

Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 4.3 + 1.13 * (6.8)
Cost of equity% = 11.98
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =20x2
1090 =∑ [(6.7*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^20x2
                   k=1
YTM = 5.9259322402
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.9259322402*(1-0.34)
= 3.911115278532
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 5.6/(103)*100
=5.44
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=3.91*0.2663+11.98*0.6754+5.44*0.0584
WACC =9.45%